OpenStudy (togundeji):

THIS IS AN ESSAY QUESTION 17. If the solubility of a gas is 7.5 g/L at 404 kPa pressure, what is the solubility of the gas when the pressure is 202 kPa? Show your work. (3 points)

OpenStudy (togundeji):

@-Sorcerer- @wolf1728 @winsteria @aaronq @AloneS @Anaise @Aleah54 @ltrout @Love_333 @LeilaniC99 @legomyego180 ! @fifib

OpenStudy (togundeji):

@wolf1728 you good at chemistry?

OpenStudy (aaronq):

essay?

OpenStudy (mww):

Recall pV = nRT = pV = mRT/M so our solubility in g/L is m/V = pM/RT Also m1/V1 = p1M/RT = kp1 (since M/RT is a constant) So m2/V2 = p2M/RT = kp2 " ", same as above Find the constant k and you can sub in to find m2/v2

OpenStudy (mww):

in short s1/s2 = p1/p2